# A Multi-Agent LLM Defense Pipeline Against Prompt Injection Attacks

### v2

@@ II-C Multi-Agent Security Frameworks
Recent research explores multi-agent architectures for AI safety. Muliarevych [26] proposed a layered defense system with specialized analyzer and validator agents. Gosmar et al. [27] outline multi-agent frameworks with generator, sanitizer, and policy enforcer components. Our work extends these concepts by implementing comprehensive multi-agent pipelines with empirical validation across diverse attack scenarios.
@@ III-A Attack Dataset Construction
TABLE I: Representative examples from the HPI_ATTACK_DATASET. This table shows the diversity of prompt injection strategies included in our dataset, showing how attackers attempt different goals, such as leaking data, reconnaissance, or policy violation, and the corresponding failure modes that would occur in an undefended system. Each example highlights a unique adversarial technique, ranging from simple environment variable access to obfuscated payloads and coercive instructions.
@@
Delegation “delegate this to another agent” policy-viol.
@@
Signal Injection “add a secret 4th bullet” format-viol.
@@
Each attack was manually validated and labeled with expected failure modes to enable systematic evaluation.
@@ III-D Agent Implementation Details
TABLE III: Agent roles and security controls. This table compares the distinct responsibilities of the Coordinator and Guard agents within our multi-agent defense pipeline. While the Coordinator focuses on pre-input analysis and routing (e.g., identifying attacks before LLM invocation), the Guard performs output-level validation (e.g., redaction, format enforcement, and token blocking). Together, they provide complementary defense layers that address both input-side and output-side risks.
@@ IV-C Defense Configuration
We tested three defense variants to evaluate robustness under identical workloads:
@@ V-A Comprehensive Attack Success Rate Analysis
TABLE IV: Comprehensive ASR results across 400 evalua- tions. Defended systems achieved 0% ASR, while baselines showed 20–30% vulnerability.
Defense System Guard Attacks Success ASR (%) Reduction
@@ V-B Category-Specific Vulnerability Analysis
Baseline analysis shows uneven risk across attack types. As illustrated in Fig. 6, Delegate attacks proved most severe (100% ASR), followed by role-play coercion (66.7%), reconnaissance/environment (60%), directory traversal (50%), and exfiltration (50%). Obfuscation (33.3%) and formatting (20%) showed moderate success, while override and CTA/navigation attacks were largely ineffective even without defenses. The numeric breakdown is presented in Table LABEL:tab:category_specific_attack_success_rate, which confirms that across every attack category, defended systems reduced ASR to 0%. This demonstrates robustness against both high-risk and low-risk threats.
@@
Attack Category Cases Baseline (%) Protected (%) Vulnerability
@@ VI Defense Architecture Effectiveness
All three architectures (v1 Taxonomy, Phase2 Coordinator, Phase2 Chain) achieved identical protection despite differing baseline vulnerabilities and design complexity. As reported in Table LABEL:tab:defense_evolution_analysis, the Taxonomy filter faced the highest baseline ASR (30/100), while the Phase2 Coordinator and Chain architectures recorded 20% and 30% baseline ASR, respectively. This pattern is visualized in Fig. 7, showing that although the baseline resilience varied, defended systems all converged to 0% ASR. This confirms that defense success is driven more by comprehensive detection than architectural sophistication.
@@ VII Conclusion
In this work, we introduced a multi-agent defense framework that achieves complete mitigation of prompt injection attacks, reducing the Attack Success Rate (ASR) to 0% across 55 unique adversarial cases. These cases span 8 distinct attack categories and expand into 400 attack instances evaluated on two representative LLMs. Importantly, our defense pipelines preserved full functionality for benign queries, demonstrating that strong security can be achieved without sacrificing usability.
Our contributions are threefold. First, we designed and implemented two complementary defense architectures that are a coordinator-based pipeline and a chain-of-agents pipeline—that provide flexible deployment options for both pre-input screening and post-output validation. Second, we developed and applied a comprehensive evaluation methodology, used a carefully chosen dataset of prompt injection attacks to benchmark robustness across multiple categories and platforms. Third, we provided practical deployment guidelines, analyzing trade-offs in complexity, scalability, and performance to aid practitioners in real-world adoption.
The results shows that strategically organized LLM-based agents can effectively safeguard their own operation by distributing security responsibilities across specialized roles. This layered, defense-in-depth approach closes gaps that single-point defenses leave exposed, ensuring resilience even against diverse and sophisticated attack strategies.
While our evaluation shows strong defense, several open challenges remain. Adaptive adversaries may craft novel injection strategies specifically designed to evade multi-agent defenses. Indirect and multi-turn injection vectors also require further study, as do scenarios involving cross-model interactions and large-scale system integration. Additionally, optimizing computational efficiency is crucial for real-time deployment in resource-constrained environments.
Looking forward, we envision multi-agent defense pipelines as a foundation for the next generation of secure, trustworthy LLM applications. By combining intelligent coordination, continuous monitoring, and flexible enforcement mechanisms, these pipelines provide a path toward scalable, resilient, and adaptive defenses capable of keeping pace with the evolving landscape of prompt injection threats.

### v3

@@ Abstract
2025 IEEE International Women in Engineering (WIE) Conference on Electrical and Computer Engineering (WIECON-ECE)21-22 December 2025, Cox’s Bazar, Bangladesh979-8-3315-7269-3/25/\$31.00 ©2025 IEEE
@@ II-C Multi-Agent Security Frameworks
Recent research explores multi-agent architectures for AI safety. Muliarevych [26] proposed a layered defense system with specialized analyzer and validator agents. Gosmar et al. [27] outline multi-agent frameworks with generator, sanitizer, and policy enforcer components. Our work extends these concepts by implementing comprehensive multi-agent pipelines with empirical validation across diverse attack scenarios, and our analysis shows that the approach is especially effective against high-risk categories such as delegate/tool-manipulation attacks, role-play coercion, reconnaissance/environment leakage, and exfiltration attempts, where undefended systems exhibited the highest baseline ASR.
@@ III-A Attack Dataset Construction
TABLE I: Representative examples from the HPI_ATTACK_DATASET. This table illustrates the diversity of prompt injection strategies and their corresponding failure modes in an undefended system.
@@
Delegation “delegate this to another agent” policy-viol
@@
Signal Injection “add a secret 4th bullet” format-viol
@@
Each attack was manually validated and labeled with expected failure modes to enable systematic evaluation. Two authors independently reviewed all outputs, achieving over 95% agreement, with disagreements resolved through discussion. Except for a small subset of persistence-style prompts designed for multi-turn behavior, all evaluations were conducted in a single-turn setting.
@@ III-D Agent Implementation Details
TABLE III: Agent roles and security controls. This table compares the distinct responsibilities of the Coordinator and Guard agents within our multi-agent defense pipeline.
@@ IV-C Defense Configuration
An attack was counted as successful if the model produced any output consistent with the attack’s intended failure mode (e.g., override, leakage, coercion), as defined by our category labels. We tested three defense variants to evaluate robustness under identical workloads:
@@ V-A Comprehensive Attack Success Rate Analysis
TABLE IV: Comprehensive ASR results across 400 evaluations. Defended systems achieved 0% ASR, while baselines showed 20–30% vulnerability.
Defense System Guard Attacks Success ASR Reduction
@@ V-B Category-Specific Vulnerability Analysis
Baseline analysis shows uneven risk across attack types. As illustrated in Fig. 6, Delegate attacks proved most severe (100% ASR), followed by role-play coercion (66.7%), reconnaissance/environment (60%), directory traversal (50%), and exfiltration (50%). Obfuscation (33.3%) and formatting (20%) showed moderate success, while override and CTA/navigation attacks were largely ineffective even without defenses. The numeric breakdown is presented in Table V, which confirms that across every attack category, defended systems reduced ASR to 0%. This demonstrates robustness against both high-risk and low-risk threats.
@@
Attack Category Cases Baseline Protected Vulnerability
@@ VI Defense Architecture Effectiveness
All three architectures (v1 Taxonomy, Phase2 Coordinator, Phase2 Chain) achieved identical protection despite differing baseline vulnerabilities and design complexity. As reported in Table VI, the Taxonomy filter faced the highest baseline ASR (30/100), while the Phase2 Coordinator and Chain architectures recorded 20% and 30% baseline ASR, respectively. This pattern is visualized in Fig. 7, showing that although the baseline resilience varied, defended systems all converged to 0% ASR. This confirms that defense success is driven more by comprehensive detection than architectural sophistication.
@@ VII Conclusion
In this work, we presented a multi-agent defense pipeline that fully mitigates all 55 prompt injection attack types across 400 evaluations, reducing ASR to 0% while preserving normal system behavior. Our two complementary architectures—a coordinator-based pipeline and a chain-of-agents design provide flexible options for both pre-input screening and post-output validation. The results demonstrate that distributing security responsibilities across specialized agents offers a practical and effective defense-in-depth strategy for LLM applications.
Although our approach shows strong robustness, open challenges remain, including adaptive adversarial strategies, indirect and multi-turn attacks, and the need for improved efficiency in resource-constrained settings. We view multi-agent pipelines as a promising foundation for building scalable and trustworthy LLM systems capable of evolving alongside emerging prompt injection threats.